train of mass 800metrictonnes is moving with a velocity 72kmph if the resistive force acting on the train is 10newtons per metrictonne find the power of engine?
1
Expert's answer
2018-06-22T11:20:08-0400
The power of engine
Power=(work done)/time
P=W/t=Fs/t=Fv
=800×10×72×(1000 )/(3600 )=160 000 W=160 kW